Give complete and combines structural formula for the following.

1) 3- Heptene
2) Butene
3) Pentene
4) 2- Hexene

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer:

1) ch3-ch2-ch=ch-ch2-ch2-ch3

2) ch2=ch-ch2-ch2

3)ch2=ch-ch2-ch2-ch3

4) ch3-ch=ch-ch2-ch2-ch3

Thanks